# Twists of Superconformal Algebras

### v2

@@ 2.3. The Question of Classifying Twists
Topological field theories have had an enormous impact in both theoretical physics and mathematics, and the most important examples arise from supersymmetric theories by a twisting procedure, introduced by Witten [Witten_1988]. We recall here a less sophisticated version of that procedure before we turn to superconformal field theories., In between, we provide a quick survey of mathematical work on these notions, mostly growing out of the pioneering but unpublished paper of Duflo and Serganova [DufloSerganova], as it provides some important structural results.
@@ 2.3.1. How Twists Appear in the Supersymmetric Case
A translation twist is an odd element $\mathcal{Q}$ such that $[\mathcal{Q},\mathcal{Q}]=0$.
@@
Recently, there has been a systematic classification of the space of twists (at least for complexified supertranslation Lie algebras), motivated by the desire to organize and analyze the behavior of twisted supersymmetric field theories. For details of this classification, see work of the first author with Safronov and Williams [ElliottSafronov, ElliottSafronovWilliams] and of Eager, Saberi and Walcher [EagerSaberiWalcher] .
@@ 2.3.2. Associated Varieties of Super Lie Algebras
Let $G_{0}$ denote the simply-connected, connected algebraic group with Lie algebra $\mathfrak{g}_{0}$. It acts on $\mathfrak{g}_{1}$ by the adjoint action and $X_{\mathfrak{g}}$ is preserved by this action. It is natural to study the stack quotient $[X_{\mathfrak{g}}/G_{0}]$, and in the case that $\mathfrak{g}$ is a supertranslation algebra, it maps to the moduli stack of twists introduced in the preceding section. (That moduli space is a further quotient, taking into account the natural isometries, like rotation, of super-Minkowski space.)
@@
  • if any pair of odd roots $\alpha,\beta\in\Delta_{1}$ satisfies ±$\pm\alpha,\pm\beta\in A$, then $\alpha+\beta\not\in\Delta_{0}$.
@@
In [GHSS] they push further and for such classical super Lie algebras, they classify the stabilizers and normalizers for each point $x\in X_{\mathfrak{g}}$, as well as the dimensions of the orbits.
@@ 2.3.3. How Twists Appear in the Conformal Case
A conformal twist is an odd element $\mathcal{Q}\in\Pi S$ such that $[\mathcal{Q},\mathcal{Q}]=\Gamma(\mathcal{Q},\mathcal{Q})=0$.
@@
Following [ElliottSafronov, EagerSaberiWalcher], we phrase our problem in the following way. Note that for the rest of this section we will discuss nilpotent elements of the complexified superconformal algebra; we will discuss the role of real forms shortly.
@@
  • Let ${\rm SConf}(p,q|S)_{0}$ is the simply-connected, connected algebraic group over $\mathbb{C}$ whose Lie algebra is $\mathfrak{sconf}(p,q|S)$. It acts on the associated variety, and if two twists are in the same orbit, then the associated dg Lie algebras are isomorphic. Thus, it is natural to consider the quotient stack
    $[X_{{\mathfrak{sconf}(p,q|S)}}/{\rm SConf}(p,q|S)_{0}],$
@@
Question 2.26.
Are there orbits among the conformal twists that do not come from the translation twists? These would leave to novel twisted field theories.
In light of List 2.12 and Theorem 2.21, Duflo and Serganova have given essentially a complete treatment of the maximal quotient stacks: they showed that for (nearly) every conformal nilpotence variety, its ${\rm SConf}(p,q|S)_{0}$-orbits consist of a finite set of points. That result implies each maximal quotient stack is a disjoint union of classifying stacks $\ast/{\rm Stab}(x)$ where $x\in X_{\mathfrak{sconf}(p,q|S)}$ is a point on the associated orbit and where ${\rm Stab}(x)$ denotes its stabilizer inside ${\rm SConf}(p,q|S)_{0}$.
@@
But for physical questions, it is also important to examine the quotients by real groups. For example, the work of [Beem_2015] examines a complex twist but working in a fixed signature, and hence with an associated real form.
@@
In other words, we raise Question 2.22 here specifically for the real forms of the superconformal algebras. In this situation, we obtain Lie subalgebras of vector fields on the conformal manifold $C(\mathbb{R}^{p,q})$, namely $\mathfrak{z}_{\mathcal{Q}}=\mathrm{Ker}(\mathcal{Q})$ and $\mathfrak{b}_{\mathcal{Q}}=\mathrm{Im}(\mathcal{Q})$. These foliations have rich geometric content, and unpacking their geometry would undoubtedly be fruitful.
This third situation also has clear physical motivation: an action functional of a genuine physical theory should be real-valued, and a Hamiltonian should be self-adjoint. A superconformal field theory has an action of a real superconformal group, so it is natural to look for twists that determine deformations as a genuine theory. That pushes us to look at the real twists.
@@ 3. The Observables of a Twisted Superconformal Theory
Given a superconformal theory with $\mathfrak{sconf}(p,q|S)$ as a symmetry algebra, a nilpotent supercharge $\mathcal{Q}$ can be added to the action as a BRST operator leading to a new theory that has the $\mathcal{Q}$-cohomology of $\mathfrak{sconf}(p,q|S)$ as a symmetry algebra. For supersymmetric theories, this process (often dubbed “twisting”) is well-known and produces new theories with remarkable properties, such as the topological field theories encoding Donaldson invariants of 4-manifolds [Witten_1988]. A crucial step is to understand the $\mathcal{Q}$-cohomology of $\mathfrak{siso}(p,q|S)$ and its geometric meaning. For superconformal theories, the twists can have more exotic geometric consequences and the associated twisted theories have surprising behavior, including the localization discovered by Beem et al [Beem_2015].
@@ 3.1. Subalgebras of $\mathcal{Q}$-Closed and $\mathcal{Q}$-Exact Elements
Let $\mathcal{Q}$ be an odd square-zero element of $\mathfrak{sconf}(p,q|S)$. As we hinted above, the properties of the $\mathcal{Q}$-twist of a superconformal field theory defined on $C(\mathbb{R}^{p,q})$ – e.g., whether the twisted theory becomes topological or holomorphic upon restriction to certain affine subspaces – are determined by the cohomology of $\mathfrak{sconf}(p,q|S)$ under the differential $[\mathcal{Q},-]$. We will make this idea more precise as follows.
@@
$[\mathcal{Q},-]\colon\mathfrak{sconf}(n|S,\mathbb{C})_{1}\to\mathfrak{sconf}(n|S,\mathbb{C})_{0},$
@@
As we discussed in Section 2.3.3 we will be particularly interested in the complexified Lie algebra (in the odd part of which the superconformal nilpotence variety is embedded), as well as real forms of the superconformal group (acting on the nilpotence variety. We define $Z_{\mathcal{Q}}$ intrinsically as a subgroup of the real superconformal group. On the other hand, for the image $B_{\mathcal{Q}}$ our must start with the (complex) Lie algebra, so we are forced to make our definition starting with the complex subgroup $B_{\mathcal{Q}}^{\mathbb{C}}$ and then intersecting with a chosen real form.
@@
$[\mathcal{Q},[X,\mathcal{Q}^{\prime}]]=[[\mathcal{Q},X],\mathcal{Q}^{\prime}]+[X,[\mathcal{Q},\mathcal{Q}^{\prime}]]=[X,Y]$
@@ 3.2.1. The Maximal Case
In the twisted theory, we focus on $\mathcal{Q}$-closed operators (i.e., we only consider an operator $\mathcal{O}$ satisfying $\mathcal{Q}\mathcal{O}=0$). Moreover, $\mathcal{Q}$-exact operators (i.e., $\mathcal{O}=\mathcal{Q}\mathcal{O}^{\prime}$ for some operator $\mathcal{O}^{\prime}$) have trivial expected value. This means that in the twisted theory we reduce to studying cohomology classes of operators $[\mathcal{O}]_{\mathcal{Q}}$: that is, two operators differing by a $\mathcal{Q}$-exact operator, i.e. $\mathcal{O}$ and $\mathcal{O}^{\prime}=\mathcal{O}+\mathcal{Q}\mathcal{O}^{\prime\prime}$ are identified in the twisted theory, $[\mathcal{O}]_{\mathcal{Q}}\simeq[\mathcal{O}^{\prime}]_{\mathcal{Q}}$. These properties have strong consequences for how the symmetries act on operators. Recall that in a topological twist of a supersymmetric theory, an $n$-point function
@@
is independent of the insertion points $\{x_{i}\}$ because translations are $\mathcal{Q}$-exact. The same argument implies that for a conformal twist, the operator product is constant as we move an insertion point along a $B_{\mathcal{Q}}$-orbit:
@@
On the other hand, if we translate just one insertion, the expected value changes: $\langle\mathcal{O}_{1}(z(x_{1}))\mathcal{O}_{2}(x_{2})\rangle$ varies with $z$, much as a 2-point function typically depends on the relative position of the insertion.
To summarize, we should analyze $C(\mathbb{R}^{p,q})$ (and configurations of multiple points in it) into $Z_{\mathcal{Q}}$-orbits and, further, into $B_{\mathcal{Q}}$-orbits. The twisted theory $\mathcal{T}^{\mathcal{Q}}$ will behave like it is “topological” along $B_{\mathcal{Q}}$-orbits and equivariant along $Z_{\mathcal{Q}}$-orbits.
@@ 3.3.1. Vertex Algebras
We begin with the structure that originally motivated the authors of [Beem_2015] to study the observables in a conformal twist, that of a vertex algebra. This structure occurs when there is a two-dimensional affine subspace, which we identify with a copy of the complex plane $\mathbb{C}\hookrightarrow C(\mathbb{R}^{p,q})$ with holomorphic coordinate $z$. We require specific conditions on the twisting supercharge $\mathcal{Q}$ that we can describe as follows:
@@
In the introduction to [Beem_2015c], it is observed that one way to obtain the desired structure of a vertex algebra in the twist is by choosing an embedding of $\mathfrak{su}(1,1|2)$ into the superconformal algebra, so that $\mathfrak{su}(1,1)$ acts on an affine subspace $\mathbb{C}$ by antiholomorphic Möbius transformations. From this point of view, the supercharge $\mathcal{Q}$ is identified with an odd element of the complexification $\mathfrak{sl}(2|2,\mathbb{C})$, which is chiral of rank $(2,0)$, and the subalgebra of $\mathcal{Q}$-exact elements is a copy of $\mathfrak{sl}(2,\mathbb{C})$ embedded diagonally in the even part of $\mathfrak{sl}(2|2,\mathbb{C})$. The authors of loc. cit. observe that such subalgebras only occur in a small number of examples of superconformal algebras (at least in Lorentzian signature), occurring in dimension 2, 4 and 6.
@@
Given a vertex algebra of local operators along an affine subspace $\mathbb{C}$ in such a twisted theory, [Argyres:2022npi] enhance the vertex algebra to a vertex algebra that includes extended operators in the directions transverse to the chosen complex plane (though not extended within the complex plane). We do not prove general results about such algebras of extended operators in this paper.

### v3

@@ 2.3. The Question of Classifying Twists
Topological field theories have had an enormous impact in both theoretical physics and mathematics, and the most important examples arise from supersymmetric theories by a twisting procedure, introduced by Witten [45]. We recall here a less sophisticated version of that procedure before we turn to superconformal field theories. In between, we provide a quick survey of mathematical work on these notions, mostly growing out of the pioneering but unpublished paper of Duflo and Serganova [20], as it provides some important structural results.
@@ 2.3.1. How Twists Appear in the Supersymmetric Case
A translation twist is an odd element $\mathcal{Q}$ such that $\mathcal{Q},\mathcal{Q}]=0$.
@@
Recently, there has been a systematic classification of the space of twists (at least for complexified supertranslation Lie algebras), motivated by the desire to organize and analyze the behavior of twisted supersymmetric field theories. For details of this classification, see [23, 22, 21] .
@@ 2.3.2. Associated Varieties of Super Lie Algebras
Let $G_{0}$ denote the simply-connected, connected algebraic group with Lie algebra $\mathfrak{g}_{0}$. It acts on $\mathfrak{g}_{1}$ by the adjoint action and $X_{\mathfrak{g}}$ is preserved by this action. It is natural to study the stack quotient $X_{\mathfrak{g}}/G_{0}]$, and in the case that $\mathfrak{g}$ is a supertranslation algebra, it maps to the moduli stack of twists introduced in the preceding section. (That moduli space is a further quotient, taking into account the natural isometries, like rotation, of super-Minkowski space.)
@@
  • if any pair of odd roots $\alpha,\beta\in\Delta_{1}$ satisfies $\pm\alpha,\pm\beta\in A$, then $\alpha+\beta\not\in\Delta_{0}$.
@@
In [27] they push further and for such classical super Lie algebras, they classify the stabilizers and normalizers for each point $x\in X_{\mathfrak{g}}$, as well as the dimensions of the orbits.
@@ 2.3.3. How Twists Appear in the Conformal Case
A conformal twist is an odd element $\mathcal{Q}\in\Pi S$ such that $\mathcal{Q},\mathcal{Q}]=\Gamma(\mathcal{Q},\mathcal{Q})=0$.
@@
Following [23, 21], we phrase our problem in the following way. Note that for the rest of this section we will discuss nilpotent elements of the complexified superconformal algebra; we will discuss the role of real forms shortly.
@@
  • Let ${\rm SConf}(p,q|S)_{0}$ is the simply-connected, connected algebraic group over $\mathbb{C}$ whose Lie algebra is $\mathfrak{sconf}(p,q|S)_{0}$. It acts on the associated variety, and if two twists are in the same orbit, then the associated dg Lie algebras are isomorphic. Thus, it is natural to consider the quotient stack
    $X_{{\mathfrak{sconf}(p,q|S)}}/{\rm SConf}(p,q|S)_{0}],$
@@
Question 2.28.
Are there orbits among the conformal twists that do not come from the translation twists? These would lead to novel twisted field theories.
In light of List 2.13 and Theorem 2.23, Duflo and Serganova have given essentially a complete treatment of the maximal quotient stacks: they showed that for (nearly) every conformal nilpotence variety, its ${\rm SConf}(p,q|S)_{0}$-orbits consist of a finite set of points. That result implies each maximal quotient stack is a disjoint union of classifying stacks $\ast/{\rm Stab}(x)$ where $x\in X_{\mathfrak{sconf}(p,q|S)}$ is a point on the associated orbit and where ${\rm Stab}(x)$ denotes its stabilizer inside ${\rm SConf}(p,q|S)_{0}$.
@@
But for physical questions, it is also important to examine the quotients by real groups. Because the group is smaller, the quotient space is a much richer object: there are many more orbits. For example, the work of [6] examines a complex twist but working in a fixed signature, and hence with an associated real form. We revisit that example in Section 5.3.1.
@@
In other words, we raise Question 2.24 here specifically for the real forms of the superconformal algebras. In this situation, we obtain Lie subalgebras of vector fields on the conformal manifold $C(\mathbb{R}^{p,q})$, namely $\mathfrak{z}_{\mathcal{Q}}=\mathrm{Ker}(\mathcal{Q})$ and $\mathfrak{b}_{\mathcal{Q}}=\mathrm{Im}(\mathcal{Q})$. These foliations have rich geometric content, and unpacking their geometry would undoubtedly be fruitful.
This third situation also has clear physical motivation: an action functional of a genuine physical theory should be real-valued, and a Hamiltonian should be self-adjoint. A superconformal field theory has an action of a real superconformal group, so it is natural to look for twists that determine deformations as a genuine theory. That pushes us to look at the real twists. We discuss this further in Section 3.3.4.
@@ 3. The Observables of a Twisted Superconformal Theory
Given a superconformal theory with $\mathfrak{sconf}(p,q|S)$ as a symmetry algebra, a nilpotent supercharge $\mathcal{Q}$ can be added to the action as a BRST operator leading to a new theory that has the $\mathcal{Q}$-cohomology of $\mathfrak{sconf}(p,q|S)$ as a symmetry algebra. For supersymmetric theories, this process (often dubbed “twisting”) is well-known and produces new theories with remarkable properties, such as the topological field theories encoding Donaldson invariants of 4-manifolds [45]. A crucial step is to understand the $\mathcal{Q}$-cohomology of $\mathfrak{siso}(p,q|S)$ and its geometric meaning. For superconformal theories, the twists can have more exotic geometric consequences and the associated twisted theories have surprising behavior, including the localization discovered by Beem et al [6].
@@ 3.1. Subalgebras of $\mathcal{Q}$-Closed and $\mathcal{Q}$-Exact Elements
Let $\mathcal{Q}$ be an odd square-zero element of $\mathfrak{sconf}(p,q|S)$. As we hinted above, the properties of the $\mathcal{Q}$-twist of a superconformal field theory defined on $C(\mathbb{R}^{p,q})$ – e.g., whether the twisted theory becomes topological or holomorphic upon restriction to certain affine subspaces – are determined by the cohomology of $\mathfrak{sconf}(p,q|S)$ under the differential $\mathcal{Q},-]$. We will make this idea more precise as follows.
@@
$\mathcal{Q},-]\colon\mathfrak{sconf}(n|S,\mathbb{C})_{1}\to\mathfrak{sconf}(n|S,\mathbb{C})_{0},$
@@
As we discussed in Section 2.3.3 we will be particularly interested in the complexified Lie algebra (the superconformal nilpotence variety is embedded in its the odd component), as well as real forms of the superconformal group (acting on the nilpotence variety). We define $Z_{\mathcal{Q}}$ intrinsically as a subgroup of the real superconformal group. On the other hand, for the image $B_{\mathcal{Q}}$ our must start with the (complex) Lie algebra, so we are forced to make our definition starting with the complex subgroup $B_{\mathcal{Q}}^{\mathbb{C}}$ and then intersecting with a chosen real form.
@@
$\mathcal{Q},[X,\mathcal{Q}^{\prime}]]=[[\mathcal{Q},X],\mathcal{Q}^{\prime}]+[X,[\mathcal{Q},\mathcal{Q}^{\prime}]]=[X,Y]$
@@ 3.2.1. The Maximal Case
In the twisted theory, we focus only on the $\mathcal{Q}$-closed operators (i.e., we only care about operators $\mathcal{O}$ such that $\mathcal{Q}\mathcal{O}=0$). Moreover, a $\mathcal{Q}$-exact operator $\mathcal{O}$ (i.e., $\mathcal{O}=\mathcal{Q}\mathcal{O}^{\prime}$ for some operator $\mathcal{O}^{\prime}$) has trivial expected value. Note, in particular, that if $\mathcal{O}_{1}$ is exact with $\mathcal{Q}\mathcal{O}^{\prime}_{1}=\mathcal{O}_{1}$, its product with any $\mathcal{Q}$-closed operators $\mathcal{O}_{1}\mathcal{O}_{2}\cdots\mathcal{O}_{n}$ is also exact:
$\mathcal{Q}(\mathcal{O}^{\prime}_{1}\mathcal{O}_{2}\cdots\mathcal{O}_{n})=\mathcal{O}_{1}\mathcal{O}_{2}\cdots\mathcal{O}_{n},$
so that such products also have vanishing expected value.
These features of the twisted theory mean we should study cohomology classes of operators $\mathcal{O}]_{\mathcal{Q}}$: that is, if $\widetilde{\mathcal{O}}=\mathcal{O}+\mathcal{Q}\mathcal{O}^{\prime}$ (i.e., the operators differ by a $\mathcal{Q}$-exact operator), then they are identified in the twisted theory $\mathcal{O}]_{\mathcal{Q}}=[\widetilde{\mathcal{O}}]_{\mathcal{Q}}$. These properties have strong consequences for how the symmetries act on operators. For instance, in a topological twist of a supersymmetric theory, an $n$-point function
@@
is independent of the insertion points $\{x_{i}\}$ because translations are $\mathcal{Q}$-exact.
The same argument implies that for a conformal twist, the operator product is unchanged if we move an insertion point along a $B_{\mathcal{Q}}$-orbit:
@@
On the other hand, if we translate just one insertion, the expected value changes: if we vary $z$, the expected value $\langle\mathcal{O}_{1}(z(x_{1}))\mathcal{O}_{2}(x_{2})\rangle$ is an interesting function in $z$, much as a 2-point function typically depends on the relative position of the insertion.
To summarize, we should analyze $C(\mathbb{R}^{p,q})$ (and configurations of multiple points in it) into $Z_{\mathcal{Q}}$-orbits and, further, into $B_{\mathcal{Q}}$-orbits. The twisted theory $\mathcal{T}^{\mathcal{Q}}$ will be “constant” along $B_{\mathcal{Q}}$-orbits and hence behave like it is “topological” in those directions. It will be equivariant along $Z_{\mathcal{Q}}$-orbits.
@@
Remark 3.6.
While the above discussion has emphasized the role of the real groups $B_{\mathcal{Q}}$ and $Z_{\mathcal{Q}}$ (and therefore the real Lie algebras $\mathfrak{b}_{\mathcal{Q}}$ and $\mathfrak{z}_{\mathcal{Q}}$ and the vector fields they generate), the complex Lie algebras $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}$ and $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}$ are frequently substantially larger. Even though these complex Lie algebras do not act geometrically on the conformal compactification, they still have a strong impact on the observables of the $\mathcal{Q}$-twisted theory. A useful analogy to bear in mind is that on $\mathbb{R}^{2}$ viewed as the complex plane, there is a complex-valued vector field $\partial/\partial\bar{z}$ that is not geometric in the sense that it does not produce integral curves, but it acts on complex-valued functions $C^{\infty}(\mathbb{R}^{2})^{\mathbb{C}}$ and the kernel of that action is the ring of holomorphic functions. We discuss this kind of phenomenon further in Section 3.3.4.
@@ 3.2.2. Working on an Affine Patch
In the next subsection we explain how to recover vertex algebras, associative algebras, and other higher algebraic structures by picking an affine patch based on the behavior of the supercharge. In particular we will choose a patch so that the closed and exact elements in the superconformal algebra include (complexified) translations of the affine patch.
We are, however, free to choose the affine patch independently from our choice of supercharge $\mathcal{Q}$, so that translations may not closed or exact.
In general the exact Lie algebra $\mathfrak{b}_{\mathcal{Q}}$ defines a foliation of the conformal compactification $C(\mathbb{R}^{p,q})$, which we may restrict to any affine patch. A $\mathcal{Q}$-twisted theory will be locally constant along the leaves of this foliation. In the examples below these leaves are affine subspaces: for other affine patches, they will instead be subspaces generated by a subgroup of the group of conformal transformations.
For example, the conformal compactification possesses an automorphism that interchanges translations and special conformal transformations in the superconformal algebra. Given a topological twist by $\mathcal{Q}$ as in Section 3.3.2, we may consider the twist by the image of $\mathcal{Q}$ under this automorphism or, equivalently, consider the $\mathcal{Q}$ twist in the “opposite” affine patch. This twist will no longer be globally topological; instead it will be locally constant along the foliation generated by the action of special conformal transformations. (Since such a foliation might be singular, there may be “defects” that appear.)
@@ 3.3.1. Vertex Algebras
We begin with the structure that originally motivated the authors of [6] to study the observables in a conformal twist, that of a vertex algebra. This structure occurs when there is a two-dimensional affine subspace, which we identify with a copy of the complex plane $\mathbb{C}\hookrightarrow C(\mathbb{R}^{p,q})$ with holomorphic coordinate $z$. We require the following specific conditions on the twisting supercharge $\mathcal{Q}$:
@@
Note something interesting about the second condition: we have an infinitesimal symmetry of the observables (which form a complex vector space) that does not generate a flow on the real manifold. We discuss this further in Section 3.3.4.
Such a twisting leads to vertex algebra.
@@
In the introduction to [9], it is observed that one way to obtain the desired structure of a vertex algebra in the twist is by choosing an embedding of $\mathfrak{su}(1,1|2)$ into the superconformal algebra, so that $\mathfrak{su}(1,1)$ acts on an affine subspace $\mathbb{C}$ by antiholomorphic Möbius transformations. From this point of view, the supercharge $\mathcal{Q}$ is identified with an odd element of the complexification $\mathfrak{sl}(2|2,\mathbb{C})$, which is chiral of rank $(2,0)$, and the subalgebra of $\mathcal{Q}$-exact elements is a copy of $\mathfrak{sl}(2,\mathbb{C})$ embedded diagonally in the even part of $\mathfrak{sl}(2|2,\mathbb{C})$. The authors of loc. cit. observe that such subalgebras only occur in a small number of examples of superconformal algebras (at least in Lorentzian signature), occurring in dimension 2, 4 and 6.
@@
Given a vertex algebra of local operators along an affine subspace $\mathbb{C}$ in such a twisted theory, [3] enhance the vertex algebra to a vertex algebra that includes extended operators in the directions transverse to the chosen complex plane (though not extended within the complex plane). In their examples, these extended operators arise by using real even elements that generate the directions. We do not prove general results about such algebras of extended operators in this paper.